4. Consider the following three shares: a. Share A is expected to provide a dividend of $10 a share forever. b. Share B is expected to pay a dividend of $5 in one year. Thereafter, dividend growth is expected to be 4% a year forever. Share C is expected to pay no dividend in year 1, $35 per share in 2 years, $50 per share in 3 years, a dividend of $1 in year 4 and then expects the dividend to grow at 3% per annum indefinitely. c. If the required rate of return for each share is 10%, which share is the most valuable? Show your workings by calculating the value for all shares. (Correct to the nearest cent.) (7 marks)
Answer #1
4)
a) Value of Share A (as per constant dividend model) = 10/10% = $     100.00
b) Value of Share B (as per constant dividend growth model) = 5/4% = $     125.00
c) Vslue of Share C is to be falculated as below:
Year Dividend PVIF at 10% PV at 10%
1 0 0.90909 $                 -  
2 35 0.82645 $          28.93
3 50 0.75131 $          37.57
4 1 0.68301 $            0.68
PV of dividends t1 to t4 = $          67.17
Continuing value of dividends = 4*1.03/(0.10-0.03) = $          58.86
PV of continuing value = 58.86*0.68301 = $          40.20
Value of share C $          99.06
ANSWER:
SHARE B IS THE MOST VALUABLE SHARE.